OpenStudy (anonymous):

Polar Polar Coordinates: Give two sets of polar coordinates that could be used to plot the given point.

OpenStudy (dumbcow):

\[r = \sqrt{x^{2}+y^{2}}\] \[\tan \theta = \frac{y}{x}\]

OpenStudy (anonymous):